Consider the reaction:
SO2Cl2(g)⇌SO2(g)+Cl2(g), Kp=2.91×103 at 298 K


In a reaction at equilibrium, the partial pressure of SO2 is 146 torr and that of Cl2 is 370 torr.

What is the partial pressure of SO2Cl2 in this mixture?
Express your answer in torr using three significant figures.

PSO2Cl2 = ? torr
1
Expert's answer
2020-06-18T14:43:10-0400

"Kp=P(SO2)* P(Cl2)\/P(SO2Cl2)"


so "P(SO2Cl2)= P(SO2)*P(O2)\/Kp"


Kp=2910

P(SO2)= 146 torr

P(Cl2) = 370 torr


"P(SO2Cl2)= 146*370\/2910"

P(SO2Cl2)= 18.564 torr